Fermi National Laboratory
 

Accelerator Update for November 29, 2006 - December 1, 2006

 

Wednesday November 29

The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring a Recycler stash (264.2E10), with the Antiproton Source (Pbar) stacking (11.5E10), with Linac, Booster, the Recycler, and the Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting on the next store.

At 10:31 AM, Operations turned off the Booster for an access to replace a modulator for an RF station (BRF6).

Operations established store 5097 at 10:48 AM with an initial luminosity of 173.92E30.

At 11:04 AM, Linac was turned off to repair an RF station (LRF5).

At 11:05 AM, technicians began repairs on the Linac Klystron Debuncher water skid.

At 12:40 PM, the Fermi Accelerators suffered a power glitch. TRF4, the Tevatron Electron Lens, Pbar, and part of Booster and Linac tripped off. The High Voltage Coordinator reported a ComEd glitch, but that no on site breakers had tripped.

Pbar resumed stacking at 2:59 PM.

At 4:34 PM, LRF5 tripped off for the first of many times.

At 5:26 PM, LRF5 tripped off and wouldn’t reset. Operators contacted an expert.

Linac resumed sending beam at 6:26 PM.

Operations reported at 6:44 PM that operators had to reset the modulator breakers for LRF5 one at a time. Operators were then able to reset the station normally.

The Recycler completed the last of three stashes of antiprotons for the evening shift at 10:53 PM.

At 11:40 PM, LRF5 tripped off for the first of two times in seven minutes. Operators had to reset modulator breakers again.

Thursday November 30

The midnight shift began with Operations monitoring store 5097 (47.24E30) and a stash (123.6E10), with Pbar stacking (28.4E10), and with NuMI and MiniBooNE taking beam.

At 4:55 AM, a pulse of beam went to the MI dump when it wasn’t expected and caused the MI beam permit to trip. The permit reset.

The Neutron Therapy Facility (NTF) reported at 7:13 AM that they would be treating three patients this morning.

The Recycler completed the last of three stashed for the midnight shift at 7:46 AM.

Operations terminated store 5097 at 9:59 AM.

Operations prepared D0 for an access at 10:18 AM.

D0 came out of access and Operations put the TeV through a dry squeeze at 11:48 AM.

The Recycler completed a stash around noon.

Operations put the TeV into shot setup at 12:34 PM.

Operations established store 5099 at 2:26 PM with an initial luminosity of 186.8E30.

MI experts began a MI-8 study and a slipstacking study at 3:10 PM.

The Recycler completed the second and last stash for the evening shift around 9:30 PM.

Friday December 1

The midnight shift began with Operations monitoring store 5099 (65.7E30), with the Recycler and Pbar stashing, and with NuMI and MiniBooNE taking beam.

The Recycler completed a stash at 12:19 AM.

The Recycler completed its last stash for the midnight shift at 5:41 AM.

At 6:49 AM, Operations reported a problem with Recycler electron cooling due to VME problems. A Pelletron expert and operators accessed MI-31 to power cycle the Pelletron. Normal cooling resumed at 7:34 AM.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow MI experts to conduct studies today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/06/2006    email Fermilab